OpenStudy (anonymous):

In triangle ABC, what is the length of side a if A = 22°, C = 115°, and c = 18? 7.44 6.74 5.62 3.44 2.18

hartnn (hartnn):

you know about sine rule ?

OpenStudy (anonymous):

no?

hartnn (hartnn):

\(\Large \dfrac{a }{\sin A } = \dfrac{c}{\sin C}\) just plug in values :)

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/trig-sine-law.html